not sure what topic is right.i am 49 yrsold been 30 yrs since school.can you show me how to work this problem and tell me why to work it that way.
answer this where do these lines meet?
x+4y=16
3x+5y=20

Answer by Alan3354(69443)

answer this where do these lines meet?
x+4y=16
3x+5y=20
------------
They meet at the point where x and y fit both equations.
To find that point using math:
--------
There's more than one way to do this. Elimination is one method.
Multiply one or both eqns to get the same coefficient for either x or y.
Multiplying the 1st eqn by 3 will make the coefficient of x 3 in both eqns.
-----------
x+4y=16 times 3 --> 3x + 12y = 48
------
3x + 12y = 48
3x + 5y = 20
--------------- Subtract
0x + 7y = 28
7y = 28
y = 4
-----
Then sub for y into any eqn
x + 4y = 16
x + 4*4 = 16
x = 16 = 16
x = 0
That's it. They cross at (0,4)